# HyperEvent: A Strong Baseline for Dynamic Link Prediction via Relative Structural Encoding

### v2

@@ IV-E Methods Comparison
IV-F Effectiveness of the Training Algorithm
TABLE IV: Efficiency of HyperEvent on Flight Dataset
@@ IV-F Effectiveness of the Training Algorithm
Training Time
per Epoch
GPU Memory MRR
@@
The efficiency gains of HyperEvent are quantified in Table IV, contrasting HyperEvent with TNCN. HyperEvent achieves a remarkable 89.83% reduction in per-epoch training time (1h 4m 33s vs. TNCN’s 10h 34m 37s) and an 84.55% reduction in GPU memory consumption (2,346MB vs. 15,189MB), while simultaneously improving prediction accuracy with a 6.95% higher MRR (87.7 vs. 82.0). These results underscore the practicality of HyperEvent as a baseline. Its efficiency is not merely an ancillary benefit but a core feature, ensuring that it can be rapidly trained and evaluated on large-scale graphs, making it an accessible and scalable reference point for the research community.
@@
IV-G Modulating Long- and Short-Term Preferences via $n_{\text{neighbor}}$ and $n_{\text{latest}}$
IV-G1 Influence of $n_{\text{neighbor}}$
@@
This divergent behavior originates from the role of $n_{\text{neighbor}}$ in defining the temporal scope for constructing event sequences, as introduced in Section III-C1. Here, $n_{\text{neighbor}}$ represents the number of most recent interaction partners stored chronologically in the adjacency table, which directly determines the historical length considered for each node. In HyperEvent, this parameter modulates the input to the relative structural encoding calculations from equations (9) to (LABEL:eq_correlation), thereby balancing long-term preferences, such as persistent node affinities over extended timelines, against short-term preferences that focus on immediate, transient interactions. Datasets like Coin and Flight, characterized by slowly evolving patterns, achieve higher MRR with larger $n_{\text{neighbor}}$ values, as they leverage richer historical context to identify event patterns. In contrast, datasets such as Review and Comment, which involve rapid, high-frequency interactions, perform better with shorter histories, as excessive context may introduce noise that obscures recent relevant events. The findings underscore the adaptability of HyperEvent in capturing dynamic graph dynamics, highlighting how tuning $n_{\text{neighbor}}$ effectively addresses the inherent temporal heterogeneities across applications.
@@
IV-G2 Influence of $n_{\text{latest}}$
@@
This behavior can be attributed to the definition of $n_{\text{latest}}$, which directly governs the number of recent events aggregated into a sequence. As elaborated in Section IV-G1, datasets primarily rely on $n_{\text{neighbor}}$ to encode long- and short-term historical preferences; $n_{\text{neighbor}}$ effectively captures node-level temporal dependencies, thereby reducing the influence of $n_{\text{latest}}$ on overall predictive performance. Furthermore, excessively increasing $n_{\text{latest}}$ escalates computational overhead due to the linear scaling of event aggregation in sequence formation, without delivering commensurate gains in accuracy for most scenarios. Consequently, adopting $n_{\text{latest}}=10$ strikes an optimal balance, providing robust predictive capabilities while maintaining computational efficiency across diverse dynamic graph environments.
IV-H Ablation Study
TABLE V: MRR of HyperEvent with/without Enhanced Correlation Vector
@@ IV-H Ablation Study
IV-H1 Without the Relative Structural Encoding Vector
This ablation evaluates the contribution of the relative structural encoding vector by comparing model performance when replacing the full 12-dimensional relative structural encoding vector (integrating equations 13, 12, and 14) with a reduced version using only 1-hop neighborhood overlap distance (Equation 13). As Table V indicates, removing multi-hop structural features causes consistent and substantial performance degradation. The MRR drops range from 4.06% on Wiki to a drastic 67.91% on Review, with all five datasets exhibiting declines exceeding 10% except Wiki. This degradation proves particularly severe for datasets like Review and Comment, where node interactions involve complex higher-order dependencies.
@@ IV-H1 Without the Relative Structural Encoding Vector
IV-H2 Without the Efficient Training Algorithm
Table VI quantitatively compares per-epoch training durations for HyperEvent with and without the proposed efficient training algorithm across five temporal graph datasets. A striking observation emerges: enabling this algorithm consistently reduces training times by over 95% across all datasets, achieving relative savings exceeding 98% on four benchmarks. Notably, the degree of acceleration correlates with the segmentation parameter $n_{\text{segment}}$. Higher $n_{\text{segment}}$ values yield more significant efficiency gains, as evidenced by the 98.07% saving on Flight ($n_{\text{segment}}=50$) versus 95.03% on Comment ($n_{\text{segment}}=20$).
This dramatic improvement stems from the algorithm’s core design: partitioning the monolithic event stream into $n_{\text{segment}}$ independent segments linearly reduces the effective event sequence length processed within each iteration from $|\mathcal{E}|$ to $\frac{|\mathcal{E}|}{n_{\text{segment}}}$. Simultaneously, tensorized implementation in Section III-D can fully leverage the massive parallel computing capabilities of GPUs while avoiding multithreading overhead. Consequently, the computational complexity per epoch scales inversely with $n_{\text{segment}}$, explaining the near-linear time reduction observed. Such efficiency is imperative for scaling HyperEvent to massive real-world temporal graphs, where raw training times without segmentation proved prohibitive (e.g., exceeding 55 hours per epoch on Flight). This approach transforms HyperEvent from a conceptual framework into a practically deployable solution for industrial-scale dynamic graphs.
TABLE VI: Per-epoch Training Time of HyperEvent with/without Efficient Training Algorithm
@@ IV-H2 Without the Efficient Training Algorithm
Acknowledgments
This should be a simple paragraph before the References to thank those individuals and institutions who have supported your work on this article.

### v3

@@ 3.5 Methods Comparison
3.6 Effectiveness of the Training Algorithm
Table 4: Efficiency of HyperEvent on Flight Dataset
@@ 3.6 Effectiveness of the Training Algorithm
Training Time per Epoch GPU Memory MRR
@@
The efficiency gains of HyperEvent are quantified in Table 4, contrasting HyperEvent with TNCN. HyperEvent achieves a remarkable 89.83% reduction in per-epoch training time (1h 4m 33s vs. TNCN’s 10h 34m 37s) and an 84.55% reduction in GPU memory consumption (2,346MB vs. 15,189MB), while simultaneously improving prediction accuracy with a 6.95% higher MRR (87.7 vs. 82.0). These results underscore the practicality of HyperEvent as a baseline. Its efficiency is not merely an ancillary benefit but a core feature, ensuring that it can be rapidly trained and evaluated on large-scale graphs, making it an accessible and scalable reference point for the research community.
@@
3.7 Modulating Long- and Short-Term Preferences via $n_{\text{neighbor}}$ and $n_{\text{latest}}$
3.7.1 Influence of $n_{\text{neighbor}}$
@@
This divergent behavior originates from the role of $n_{\text{neighbor}}$ in defining the temporal scope for constructing event sequences, as introduced in Section 2.3.1. Here, $n_{\text{neighbor}}$ represents the number of most recent interaction partners stored chronologically in the adjacency table, which directly determines the historical length considered for each node. In HyperEvent, this parameter modulates the input to the relative structural encoding calculations from equations (9) to (LABEL:eq_correlation), thereby balancing long-term preferences, such as persistent node affinities over extended timelines, against short-term preferences that focus on immediate, transient interactions. Datasets like Coin and Flight, characterized by slowly evolving patterns, achieve higher MRR with larger $n_{\text{neighbor}}$ values, as they leverage richer historical context to identify event patterns. In contrast, datasets such as Review and Comment, which involve rapid, high-frequency interactions, perform better with shorter histories, as excessive context may introduce noise that obscures recent relevant events. The findings underscore the adaptability of HyperEvent in capturing dynamic graph dynamics, highlighting how tuning $n_{\text{neighbor}}$ effectively addresses the inherent temporal heterogeneities across applications.
@@
3.7.2 Influence of $n_{\text{latest}}$
@@
This behavior can be attributed to the definition of $n_{\text{latest}}$, which directly governs the number of recent events aggregated into a sequence. As elaborated in Section 3.7.1, datasets primarily rely on $n_{\text{neighbor}}$ to encode long- and short-term historical preferences; $n_{\text{neighbor}}$ effectively captures node-level temporal dependencies, thereby reducing the influence of $n_{\text{latest}}$ on overall predictive performance. Furthermore, excessively increasing $n_{\text{latest}}$ escalates computational overhead due to the linear scaling of event aggregation in sequence formation, without delivering commensurate gains in accuracy for most scenarios. Consequently, adopting $n_{\text{latest}}=10$ strikes an optimal balance, providing robust predictive capabilities while maintaining computational efficiency across diverse dynamic graph environments.
3.8 Ablation Study
Table 5: MRR of HyperEvent with/without Enhanced Correlation Vector
@@ 3.8 Ablation Study
3.8.1 Without the Relative Structural Encoding Vector
This ablation evaluates the contribution of the relative structural encoding vector by comparing model performance when replacing the full 12-dimensional relative structural encoding vector (integrating equations 13, 12, and 14) with a reduced version using only 1-hop neighborhood overlap distance (Equation 13). As Table 5 indicates, removing multi-hop structural features causes consistent and substantial performance degradation. The MRR drops range from 4.06% on Wiki to a drastic 67.91% on Review, with all five datasets exhibiting declines exceeding 10% except Wiki. This degradation proves particularly severe for datasets like Review and Comment, where node interactions involve complex higher-order dependencies.
@@ 3.8.1 Without the Relative Structural Encoding Vector
3.8.2 Without the Efficient Training Algorithm
Table 6 quantitatively compares per-epoch training durations for HyperEvent with and without the proposed efficient training algorithm across five temporal graph datasets. A striking observation emerges: enabling this algorithm consistently reduces training times by over 95% across all datasets, achieving relative savings exceeding 98% on four benchmarks. Notably, the degree of acceleration correlates with the segmentation parameter $n_{\text{segment}}$. Higher $n_{\text{segment}}$ values yield more significant efficiency gains, as evidenced by the 98.07% saving on Flight ($n_{\text{segment}}=50$) versus 95.03% on Comment ($n_{\text{segment}}=20$).
This dramatic improvement stems from the algorithm’s core design: partitioning the monolithic event stream into $n_{\text{segment}}$ independent segments linearly reduces the effective event sequence length processed within each iteration from $|\mathcal{E}|$ to $\frac{|\mathcal{E}|}{n_{\text{segment}}}$. Simultaneously, tensorized implementation in Section 2.4 can fully leverage the massive parallel computing capabilities of GPUs while avoiding multithreading overhead. Consequently, the computational complexity per epoch scales inversely with $n_{\text{segment}}$, explaining the near-linear time reduction observed. Such efficiency is imperative for scaling HyperEvent to massive real-world temporal graphs, where raw training times without segmentation proved prohibitive (e.g., exceeding 55 hours per epoch on Flight). This approach transforms HyperEvent from a conceptual framework into a practically deployable solution for industrial-scale dynamic graphs.
Table 6: Per-epoch Training Time of HyperEvent with/without Efficient Training Algorithm